# Technological Complexity Based on Japanese Patent Data

### v2

@@ Abstract
As international competition intensifies in technologies, nations need to identify key technologies to foster innovation. However, the identification is difficult because a technology is independent, and therefore it has complex nature. Here, this study aims to assess the patent technological fields by applying Technological Complexity Index from a corporate perspective, addressing its underutilization in Japan despite its potential. By utilizing carefully processed patent data from fiscal years 1981 to 2010, we analyze the bipartite network which consists of 1,938 corporations and 35 or 124 technological fields. Our findings provide quantitative characteristics of the ubiquity and the sophistication for each patent fields, the detailed technological trends that reflect the social context, and methodological stability for policymakers and researchers, contributing to targeted innovation strategies in Japan.
@@ Introduction
In the twenty first century, the global landscape of innovation has grown markedly competitive. Nations around the world have designed diverse industrial policies to foster innovation, recognizing it as a vital factor for sustaining economic competitiveness. Despite the abundance of strategies, one persistent difficulty to identify the core technologies and actors that drive innovation remains. The difficulty arises from the fact that innovation is an intangible phenomenon emerging from an interconnected ecosystem composed of countries, cities, corporations, and technologies. Understanding this complex environment requires methodologies capable of pinpointing the central elements that contribute to innovation amid intricate interdependencies. Globally, different regions and institutions have approached this challenge from various angles. In the European Union, the concept of Smart Specialisation [1, 2] has helped identify key domains for policy support, while in Japan, science and technology strategies continue to evolve [3]. For supporting them, a range of indicators have been proposed to assess innovation across manufacturing, services, and advanced materials [4, 5]. Some studies focus on either national-level [2, 6], regional-level [7, 8, 9, 10] or corporate-level analyses [11, 12, 13]. Moreover, recent research has addressed diverse sectors such as Industry 4.0 [2], pharmaceutical biotechnology [14, 15, 16, 17], and service innovation [4], reflecting the breadth of modern innovation ecosystems. Among these approaches, a notable stream of research has emerged from the study of product complexity derived from Hidalgo-Hausumann (HH) algorithm[18]. The algorithm quantifies sophistication by applying network analysis to trade data and it has been adapted to different contexts such as economic, patents, and technological capabilities [19, 20, 21, 22, 23]. The algorithm also has proven instrumental in highlighting technologies and capabilities that hold particular importance.
Building on this foundation, Balland and Rigby [24] developed the Technological Complexity Index (TCI) to measure the complexity of technological fields using patent data. The TCI captures the nuanced interplay between technological ubiquity (i.e., how few players are capable of producing it) and sophistication (i.e., ubiquitous and produced by key players) [7]. Various extensions and complementary metrics have been proposed, including fitness-based approaches [25, 26, 27] and additional frameworks for classification and measurement [28, 29, 30, 31, 32]. More recently, the method has been extended to Asian contexts and corporate-level analyses, supported by the increasing availability of detailed patent databases [33, 34]. Despite these advancements, comprehensive applications of these methods in the Japanese context remain limited. In addition, although Japan has experienced prolonged economic stagnation, often described as the “Lost Decades”, it can hardly be denied that Japanese government cannot effectively identify the technologies that will drive renewed innovation and international competitiveness [35]. Although some studies, such as the work on automotive suppliers[13], have employed related approaches, there is a clear lack of a broad analysis that would illuminate the full landscape of Japanese technologies through the lens of complexity. This study addresses the existing gap by applying the HH algorithm to patent data from the Patent Office of a country, which encompasses a diverse range of technological fields. A bipartite network is constructed to link domestic corporations with these technological domains, and a subsequent projection into a monopartite network facilitates the derivation of TCI. Represented by the second largest eigenvector of the network adjacency matrix, TCI enables a simultaneous evaluation of both technological spillovers and scarcity. An analysis of temporal trends over the period 1981 to 2010 demonstrates that the proposed method yields a robust and detailed depiction of the technological landscape, thereby surpassing conventional quantitative measures such as simple patent counts. The assessment reveals detailed aspects of technological capabilities and their evolution over time. This study contributes to a comprehensive understanding of the innovation ecosystem by identifying the central technological fields through the lens of complexity. The insights obtained herein may inform the development of more effective innovation policies and strategies, and they establish a methodological framework that can be extended to comparative studies in other contexts. The rest of this paper is organized as follows. In Results Results section, we present the ubiquity and complexity of technological fields presented by the patent data, followed by a detailed analysis of technological trends. In the Discussion Discussion section, we provide a critical analysis of our findings and discuss the implications of our results. Finally, in the Methods Methods section, we detail the data processing and methods used in our analysis.
@@ Results
In this study, we employ the HH algorithm to derive quantitative indicators that characterize each technology (see Methods Methods). First, we define ubiquity (or degree centrality) $K_{T,0}$ as the number of corporations connected to a technology; a lower value indicates that the technology is more exclusive. Similarly, the average diversity $K_{T,1}$, which measures the variety of corporations connected to a technology, implies that a higher value reflects diversified publication, whereas a lower value suggests specialization. In many cases, the average of each indicator is used as a threshold to distinguish high from low values [18, 36].
@@
Another key indicator is the TCI, which represents the sophistication required to produce a technology. A threshold of zero divides TCI into high and low categories: a high TCI implies that a technology demands a complex combination of unique inputs and advanced technical expertise which explains why only a few advanced corporations can publish it competitively, whereas a low TCI indicates reliance on widely available capabilities, allowing many corporations to participate.
To analyze the structure of corporate patent publication as depicted by the bipartite graph, we focus on the relationships among these three indicators. The mapping of technologies is based on the Schmoch system [28] (Panel A in Fig. 1), where $K_{T,1}$ and $K_{T,0}$ are divided by their mean values, and TCI is segmented by a zero threshold.
Panel B in Fig. 1 shows that there is no correlation between $K_{T,0}$ and $K_{T,1}$ (Pearson correlation coefficient $r=0.064$), suggesting that technologies widely held by corporations do not necessarily correspond to more diversified production portfolios. In particular, the lower-left quadrant of Panel B represents sectors with both low ubiquity and low corporate diversification. For example, sectors corresponding to Electrical Engineering are predominantly located in this quadrant.
@@
On the other hand, Panel C in Fig. 1 reveals a moderate positive correlation between $K_{T,0}$ and TCI (Pearson correlation coefficient $r=0.594$). In this context, Electrical Engineering exhibits low ubiquity, low average diversity, and low TCI values, indicating that corporations in this field tend to concentrate their innovation efforts within narrowly defined domains.
The relative positioning of technologies was further confirmed using a finer classification based on International Patent Classification (IPC) classes (SI Fig. S1). Focusing on the five sectors defined by Schmoch [28], we observe that in four sectors the relationship between ubiquity and average diversity is either negligible or moderately correlated. Notably, the Electrical Engineering sector exhibits a clear positive correlation between ubiquity and TCI (Pearson correlation coefficient $r=0.738$), indicating that among corporations specializing in Electrical Engineering, some produce a comprehensive range of ubiquitous technologies while others focus on rare, specialized ones.
@@
In contrast to Electrical engineering technologies, Chemistry and Pharmaceuticals technologies mostly exhibit high TCI and the high average diversity. (Panel D in Fig. 1, Pearson correlation coeffecience $r=0.316$) These sectors are characterized by rarity and high sophistication and are typically developed by more diversified corporations. Moreover, corporations in these domains are connected not only to fields with low ubiquity but also to those that are more common (e.g., Biotechnology, Analysis of Biological Materials, and Measurement as shown in the right quadrant of Panel C in Fig. 1). A detailed examination reveals a slight negative correlation between ubiquity and average diversity (SI Fig. S1), suggesting that in sectors with high technological complexity, corporations with more comprehensive technological portfolios tend to produce rare technologies. Furthermore, the IPC classes corresponding to Biotechnology and Measurement, show high TCI values in fields related to the production of sugar and yeast (C13, A21, C12), drug discovery (A61), and the preservation and management of these technologies (C07, A23). These findings are consistent with the socio-technical context of the period, particularly reflecting the emergence of drug discovery technologies in the 1980s and 1990s [14, 15]. Moreover, the significant roles of food chemistry and biotechnology during the 1990s and 2000s underscore their contribution to the development of additives that enabled safer food preservation and processing not only in Japan but globally [16, 17].
@@
Figure 1: (A) The legend of the five classifications defined by Schmoch [28], along with an explanation of the technology codes used in (B), (C), and (D). (B) Pearson correlation between $K_{T,0}$ and $K_{T,1}$ ($r=0.064$). (C) Pearson correlation between $K_{T,0}$ and TCI ($r=0.594$). (D) Pearson correlation between TCI and $K_{T,1}$ ($r=0.316$). In (B) through (D), the black lines indicate the mean values of $K_{T,0}$ and $K_{T,1}$, and $\text{TCI}=0$ is also shown with a black line. Each dot is color-coded according to (A).
@@
Traditional regional approach measures complexity as a relative index that reflects the outcomes of agglomeration economies at the country or city level [37] and relates to spatial inequality. This measure is strongly affected by the degree of urbanization in Japan. However, the regional approach depends on the aggregation level of regions or technologies [19]. When the HH algorithm is applied to patent data, finer classifications such as IPC classes lead to inconsistent behaviour of the complexity index when compared with coarse classifications such as the Schmoch system [32]. This lack of consistency applies to the evaluation of regional complexity and also occurs in the evaluation of technology complexity owing to the symmetry of the bipartite network.
In this study we compare the TCI from the regional approach with the TCI from the corporate approach. We examine the differences in inequality and instability of complexity between regions and corporations by comparing results obtained with coarse and fine classifications (Panel A in Fig. 2).
The scaled TCI of the Schmoch system shows that in the regional approach 26 out of 35 classes fall in the range of 75 to 100. In the corporate approach, only 15 classes lie in that range, whereas the other classes display greater differences among technologies in the regional approach. This difference in heterogeneity stems from the corporate approach being able to characterize each technology in a more detailed manner. The HH algorithm defines ubiquity as the degree of a technology node and this measure depends on the number of paired nodes. The regional approach comprised roughly a few hundred region nodes, such as countries or EU cities[19, 38]. The corporate approach has corporation nodes numbering from one thousand to several ten thousand. The iterative calculation of TCI can therefore reflect a wider range of ubiquity.
@@
These consistent and detailed trends are observed not only over a 30 year period but also in evaluations over shorter periods such as 5 years. For example, in the IPC classes corresponding to Food chemistry, the trend of high TCI is consistent in the 2000s. Technologies with high scores as C13 and A21 and those that decline as A23 and A01 are captured in a consistent manner (SI Fig. S2).
The corporate approach also permits a relative evaluation of technologies in a specific region that was not possible with the regional approach. The regional approach evaluates each region only by whether it possesses a technology that has been relatively evaluated at the national level. The corporate approach can assess which technology is important within that region. For example, the relationship between the TCI of each IPC class for three prefectures with many corporations and the TCI for all corporations shows that Aichi, the prefecture where a lot of co-manufacturing companies, such as Toyota, exhibits a trend different from Tokyo and Osaka (Fig. 3). In Aichi, the technologies with a relatively high evaluation even though the overall evaluation is low are the strengths of corporations located in that region.
Figure 2: (A) The TCI scores of IPC first Class (N=124) and Schmoch (N=35) for both prefectures and corporations. The scores have been scaled to values ranging from 0 to 100. Each IPC is plotted against the corresponding Schmoch axis, and in cases of many-to-many correspondences, each pair is plotted individually. (B) The distribution of the absolute differences between TCI of IPC and of Schmoch systems.
@@ Discussion
In this study we applied the TCI to Japanese corporate patent data, targeting corporations and all patent technology fields, to identify sophisticated technological areas that are key to fostering innovation. Our findings highlighted high complexity in the Pharmaceuticals and Chemistry sectors, a result that may reflect the societal conditions of the period, often described as the dawn of the pharmaceutical industry in Japan. Importantly, the corporate level TCI preserved the same relative ranking of technologies even after switching from coarse Schmoch fields to finer IPC classes, demonstrating measurement invariance in contrast to earlier regional studies in which this ordering changed. Although previous work argued that evaluation results of these technological fields may converge because of their classification methods [19], the corporate level approach produced consistent evaluations even with finer classifications. This finding contrasts with the traditional regional approach, which often relies on coarse classifications to keep stability. The ability of the corporate approach to give consistent results at finer granularity suggests its potential to provide more accurate and detailed insight into technological trends within specific regions, industries, or other limited scopes. Limitations specific to the present method remain, particularly because capturing changes in technological trends requires both long term and short term evaluations that detect emerging fields and shifts in innovation patterns, yet such finer grained analyses introduce challenges in data processing. For instance, the calculation of the TCI is sensitive to the number of patents by corporation and technological field in each period. Shortening the aggregation period may introduce corporations with only a few patents, which adds noise to the TCI and reduces the stability of the results [32]. This limitation reflects a shortcoming of the HH algorithm, which evaluates network complexity from node degrees in a bipartite graph and therefore relaxes the conditions for link formation when the data are sparse. Several mathematical and algorithmic improvements can mitigate this issue. The Fitness Complexity algorithm [25] uses nonlinear iterative calculations that better account for network structure and shows strong correlations with indicators based on the HH algorithm while improving robustness in the evaluation of regional and corporate competitiveness [26, 27]. In addition, links can be weighted by citation counts instead of raw patent counts to obtain a more systematic and scalable measure of technological effort. Applying these enhancements within our framework would allow deeper insight into Japan’s industrial technology landscape. Future research should extend this work in several directions. Robustness checks that compare the corporate level TCI with alternative indicators such as citation counts, methods that predict temporal changes in the bipartite graph using machine learning, and cross national comparisons are promising next steps. Ultimately, this study represents a cornerstone for understanding the characteristics of domestic industries in Japan through the lens of technological complexity. By identifying key technological areas, the findings offer guidance that can inform industrial policies and strategies, especially during a period marked by rapid progress in artificial intelligence and growing international competition.

### v3

@@ Abstract
As international competition intensifies in technologies, nations need to identify key technologies to foster innovation. However, the identification is challenging due to the independent and inherently complex nature of technologies. Traditionally, regional analyses of technological portfolios have been limited to binary evaluations, indicating merely whether a region specializes in a technology, or relying on the average Technological Complexity Index (TCI) of the specialized technologies. This study proposes that evaluating TCI at the corporate level could provide finer granularity and more detailed insights. To address the underutilization of corporate-level TCI assessments in Japan, this study applies the Technological Complexity Index using carefully processed patent data spanning fiscal years 1981 to 2010. Specifically, we analyze a bipartite network composed of 1,938 corporations connected to technological fields categorized into either 35 or 124 classifications. Our findings quantitatively characterize the ubiquity and sophistication of each technological field, reveal detailed technological trends reflecting broader societal contexts, and demonstrate methodological stability even when employing finer technological classifications. Additionally, our corporate-level approach allows consistent comparisons across different regions and technological fields, clarifying regional advantages in specific technologies. This refined analytical framework offers policymakers and researchers robust, targeted insights, thereby significantly contributing to innovation strategy formulation in Japan.
@@ Introduction
In the twenty-first century, intensifying international competition has made technological capability a central element of national competitiveness. Nations seek to foster environments that support the development of sophisticated technologies, as these capabilities are fundamental to sustainable economic growth. [1, 2, 3] The ecosystems that produce such capabilities are, however, inherently complex, involving interactions among countries, corporations, and technologies [4]. This complexity makes it difficult to accurately identify the key technologies and actors within the system. Understanding this structure is crucial for developing effective science and technology strategies to strengthen a nation’s competitive advantage.
Historically, the assessment of technological capabilities has predominantly relied on quantitative metrics derived from bibliographic data of patents and publications, ranging from simple patent counts to more complex citation-based indicators, including Technology Relevance (TR) [5, 6]. While such indicators offer certain insights by considering citation practices, patent age, and technological fields, they suffer from significant shortcomings. Their inherent dependence on the long-term accumulation of citations makes them susceptible to biases from citation behaviors, patent examiners, and strategic patenting. Crucially, these metrics fail to capture the core aspect of technological complexity, namely the intricate interplay among corporations and technologies. Consequently, approaches that rely solely on these data provide a limited perspective and inadequately reflect the qualitative dimensions of technological capability.
The HH algorithm provides a powerful solution by considering both the diversity of a region and the ubiquity of its products or technologies [7]. This dual perspective allows the methodology to distinguish between components and quantify their complexity, revealing detailed economic structures not captured by traditional metrics. The utility of this foundational framework has led to its extension into several distinct streams of research[8, 9, 10, 11, 12]. One major development is the introduction of complementary metrics such as fitness-based approaches [13, 14, 15]. Another significant stream, which is the focus of this study, involves the adaptation of the algorithm to patent data, pioneered by Balland and Rigby to measure technological complexity [16]. Building on this work, recent studies have further extended the patent-based analysis to contexts such as Asian economies [17, 18]. Despite this wide range of applications, the focus has predominantly remained at a macro level, such as on cities or regions [8]. Consequently, even within the stream of patent analysis, comprehensive validation at the corporation level is still limited.
To address this gap, our study applies the HH algorithm to an extensive set of Japanese patent records to quantify technological complexity at the corporation level. By constructing bipartite networks linking corporations to technology classes, we develop a detailed measure of technological sophistication. Our analysis investigates the relationships between technological complexity, ubiquity, and the average corporate diversity of patent technologies, providing new insights into the nature of advanced capabilities. The model developed at the corporation level provides a more robust and consistent assessment of technology than regional-level models, particularly for finely classified technological fields. Importantly, this study shows that a corporation-level approach enables a direct comparison of technological characteristics across different regions. By focusing on corporations within specific locales, we can identify distinct industrial specializations and reveal which technologies are complex within each region’s unique context. The framework and insights from this study provide a powerful tool for creating tailored industrial policies and more accurately identifying the key technologies vital for future growth.
The rest of this paper is organized as follows. In Introduction section, we present the ubiquity and complexity of technological fields presented by the patent data, followed by a detailed analysis of technological trends. In the Finer Technological Trends and Consistency section, we provide a critical analysis of our findings and discuss the implications of our results. Finally, in the Discussion section, we detail the data processing and methods used in our analysis.
@@ Results
In this study, we use the HH algorithm to quantify the structural characteristics of technologies based on patent publications from 1,938 corporations filed between fiscal years 1981 and 2010 (see Discussion section). First, we define ubiquity (or degree centrality) $K_{T,0}$ as the number of corporations connected to a technology which lower value indicates that the technology is more rare. Similarly, the average diversity $K_{T,1}$, which measures the variety of corporations connected to a technology, implies that a higher value reflects diversified publication, whereas a lower value suggests specialization. Generally, the average of each indicator is used as a threshold to distinguish high from low values [19, 16, 20].
@@
Another key indicator is TCI, which represents the sophistication of a technology by reflecting how advanced the knowledge required for its production. A threshold of zero divides TCI into high and low categories: a high TCI implies that a technology demands a complex combination of unique inputs and advanced technical expertise which indicates only a few advanced corporations can publish it competitively, whereas a low TCI indicates reliance on widely available capabilities, allowing many corporations to participate.
To systematically evaluate technological characteristics, we map the relationships among ubiquity $K_{T,0}$, average diversity $K_{T,1}$, and TCI across patent technologies. (Fig. 1) Technologies are classified according to the Schmoch system [21], as illustrated in Panel A of Fig. 1. Values for ubiquity and average diversity are normalized by their respective means, while TCI is segmented by the zero threshold.
We find a negligible correlation between ubiquity $K_{T,0}$ and average diversity $K_{T,1}$ (Pearson correlation coefficient $r=0.064$), indicating that widely disseminated technologies are not necessarily associated with diversified corporate portfolios (Panel B in Fig. 1). Technologies within Electrical Engineering, for example, typically show low ubiquity and low average diversity, indicating concentrated innovation within narrowly defined technological domains.
@@
Conversely, ubiquity $K_{T,0}$ and TCI exhibit a moderate positive correlation (Pearson correlation coefficient $r=0.594$), suggesting that technologies associated with fewer corporations often require complex capabilities (Panel C in Fig.1). In this context, Electrical Engineering exhibits low ubiquity, low average diversity, and low TCI values, indicating that corporations in this field tend to concentrate their innovation efforts within narrowly defined domains.
Detailed analysis using International Patent Classification (IPC) classes further supports these observations (Fig. A1). Within the five major sectors defined by Schmoch [21], the relationship between ubiquity and average diversity is predominantly weak or moderate. Specifically, Electrical Engineering displays a strong positive correlation between ubiquity and TCI (Pearson correlation coefficient $r=0.738$), highlighting the coexistence of corporations specializing in broadly applicable technologies and those focusing on niche, specialized innovations.
@@
In contrast to Electrical engineering technologies, Chemistry and Pharmaceuticals technologies mostly exhibit high TCI and the high average diversity (Panel D in Fig. 1, Pearson correlation coeffecience $r=0.316$). These sectors are characterized by rarity and high sophistication and are typically developed by more diversified corporations. Moreover, corporations in these domains are connected not only to fields with low ubiquity but also to those that are more common (e.g., Biotechnology, Analysis of Biological Materials, and Measurement as shown in the right quadrant of Panel C in Fig. 1). A detailed examination reveals a slight negative correlation between ubiquity and average diversity (Fig. A1), suggesting that in sectors with high technological complexity, corporations with more comprehensive technological portfolios tend to produce rare technologies. Furthermore, the IPC classes corresponding to Biotechnology and Measurement, show high TCI values in fields related to the production of sugar and yeast (C13, A21, C12), drug discovery (A61), and the preservation and management of these technologies (C07, A23). These findings are consistent with the socio-technical context of the period, particularly reflecting the emergence of drug discovery technologies in the 1980s and 1990s [22, 23]. Moreover, the significant roles of food chemistry and biotechnology during the 1990s and 2000s underscore their contribution to the development of additives that enabled safer food preservation and processing not only in Japan but globally [24, 25].
@@
Figure 1: (A) The legend of the five classifications defined by Schmoch [21], along with an explanation of the technology codes used in (B), (C), and (D). (B) Pearson correlation between $K_{T,0}$ and $K_{T,1}$ ($r=0.064$). (C) Pearson correlation between $K_{T,0}$ and TCI ($r=0.594$). (D) Pearson correlation between TCI and $K_{T,1}$ ($r=0.316$). The black lines indicate the mean values of $K_{T,0}$ and $K_{T,1}$, and $\text{TCI}=0$ is also shown with a black line. Each dot is color-coded according to (A).
@@
The traditional regional approach measures technological complexity as a relative index reflecting regional technological capabilities, shaped by the economic advantages arising when corporations cluster within particular countries or cities, known as agglomeration economies [26]. Consequently, this method inherently captures spatial inequalities and is strongly influenced by urbanization levels in Japan. In terms of evaluating a region’s technological portfolio, this approach typically provides limited insights, either indicating binary specialization in technologies (0 or 1) or offering only the average TCI of the specialized technologies. Furthermore, finer classifications, such as IPC classes, tend to produce less consistent complexity indices compared to broader classifications like the Schmoch system [27]. Such inconsistencies appear in both regional and technological complexity evaluations due to the symmetric nature of the bipartite network structure underlying these calculations.
In this study, we propose examining technological complexity at the corporate level to overcome these limitations and achieve a more detailed understanding of regional technological portfolios. Specifically, by calculating TCI at the corporate level, we explore whether it is possible to capture more granular insights into technological capabilities within regions. By comparing TCI results derived from regional and corporate approaches, we assess how classification granularity influences the evaluation of inequality and stability of technological complexity across regions and corporations (Panel A in Fig. 2). In this study, we address these inconsistencies by comparing TCI results obtained from regional and corporate approaches, analyzing how classification granularity affects inequality and stability of complexity across regions and corporations (Panel A in Fig. 2).
The scaled TCI of the Schmoch system shows that in the regional approach 26 out of 35 classes fall in the range of 75 to 100. In the corporate approach, only 15 classes lie in that range, whereas the other classes display greater differences among technologies in the regional approach. This difference in stems from the corporate approach being able to characterize each technology in a more detailed manner. The HH algorithm defines ubiquity as the degree of a technology node and this measure depends on the number of paired nodes. The regional approach comprised roughly a few hundred region nodes, such as countries or EU cities[8, 28]. The corporate approach has corporation nodes numbering from one thousand to several ten thousand. The iterative calculation of TCI can therefore reflect a wider range of ubiquity.
@@
These consistent and detailed trends are observed not only over a 30 year period but also in evaluations over shorter periods such as 5 years. For example, in the IPC classes corresponding to Food chemistry, the trend of high TCI is consistent in the 2000s. Technologies with high scores as C13 and A21 and those that decline as A23 and A01 are captured in a consistent manner (SI Fig. A2).
Moreover, the corporate approach facilitates relative technological evaluations within specific regions, a capability absent in the traditional regional approach, which relies solely on the national-level presence of technologies. The corporate method identifies strategically important technologies within individual regions. For example, comparative analyses of TCI values for IPC classes among three representative prefectures, Tokyo, Osaka, and Aichi, highlight distinct local technological strengths. Aichi, notable for its concentration of major manufacturing corporations including Toyota, exhibits a technological profile markedly different from Tokyo and Osaka (Fig. 3). Technologies showing relatively higher local complexity despite lower overall national rankings underline region-specific corporate advantages.
Figure 2: (A) The TCI scores of IPC first Class (N = 124) and Schmoch (N = 35) for both prefectures and corporations. The scores have been scaled to values ranging from 0 to 100. Each IPC is plotted against the corresponding Schmoch axis, and in cases of multiple correspondences (i.e., Schmoch category “Chemical Engineering” maps to IPC classes B01–B08, C14, etc.), each pair is plotted individually. (B) The distribution of the absolute differences between TCI of IPC and of Schmoch systems.
@@ Discussion
This study applies the Technological Complexity Index (TCI) to Japanese corporate patent data across various technological fields to identify sophisticated areas critical to innovation. Accurate assessment of technological complexity is essential for policymakers and corporate strategists aiming to maintain competitive advantages amidst intensifying global competition. Previous regional approaches have struggled with instability when employing detailed technological classifications, resulting in coarse and potentially misleading insights. This research provides three primary contributions. First, it demonstrates methodological robustness at the corporate level. Unlike regional approaches, corporate-level TCI retains stable technological rankings even when classifications become highly detailed. Our results reveal notable complexity in the Pharmaceuticals and Chemistry sectors, likely reflecting Japan’s industrial focus during the pharmaceutical industry’s formative period. This consistency across classification scales represents significant progress compared to previous regional analyses that lost reliability when switching from broader to finer classifications. Second, the empirical findings underscore the advantage of evaluating technological complexity at the corporate level. Prior research suggested that complexity rankings could converge due to methodological limitations of classification systems [8]. However, our corporate-level analysis maintains clear distinctions between technological fields even at granular classification levels. This evidence highlights the corporate approach’s potential to provide deeper and more accurate insights into technological dynamics within specific industries or regions. Third, our results offer practical implications for industrial policy and strategy. The identification of high-complexity technological sectors informs targeted investment and innovation strategies, especially relevant amid rapid advances in artificial intelligence and growing global competition. Corporate strategists can leverage these insights to allocate resources effectively, fostering technological leadership and competitive differentiation. Nevertheless, our study faces several limitations. Capturing dynamic shifts in technological trends requires a balance between long-term analysis, which ensures stability, and short-term assessments, which detect emerging innovations. Shortening aggregation periods introduces challenges due to the sensitivity of TCI calculations to sparse patent data, thus decreasing result stability [27]. The HH algorithm underlying TCI calculations evaluates complexity based on node connections within a bipartite network, making it vulnerable to inaccuracies when patent data are sparse. To address these challenges, future studies could integrate advanced methods such as the Fitness Complexity algorithm [13], which employs nonlinear iterative processes to better capture network structures, enhancing robustness and reliability [14, 15]. Additionally, weighting links by citation counts rather than raw patent numbers could offer a more precise representation of technological effort and impact. Future research directions include comparative robustness analyses of TCI with alternative indicators such as patent citations, application of machine learning techniques to predict technological shifts, and cross-national comparative studies. Such explorations will deepen our understanding of technological complexity and provide richer guidance for industrial development strategies. Ultimately, this research significantly advances our understanding of technological complexity at the corporate level in Japan. By clearly identifying key technology sectors, our findings inform strategic planning and policy-making, positioning Japan advantageously in an era marked by technological acceleration and intensified international competition.